6) For each of the following utility functions derive directly from the definition not using the formula(s) from class or the tert - the MRS (marginal rate of substitution) of y for r at (2,4) (a) ua(z, y)= 1/2y1/2 (b) uo(x, y) 2y/2 (b) u(r, y)4y4 (d) udr,y)-1/ 1/y

Please explain in slow steps how MRS can be derived using definition of MRS from the ratio of partial derivatives. No specific information is needed about class formulas.
Answer #1

MRSyx = Marginal utility of x/ marginal utility of y

Marginal utility of x = du/dx

Marginal utility of y = du/dy

(a) MUx = du/dx = 1/2 (y/x)1/2

MUy = du/dy = 1/2 (x/y)1/2

MRS = y/x

MRS at (2,4) = 4/2 = 2

(b) MUx = 1/2 1/ x1/2

MUy = 1/2 1/ y1/2

MRS = (y/x )1/2 = 21/2

c) MUx = 1/ x3/4

MUy = 1/ y3/4

MRS = (y/x)3/4 = (2)3/4 = 2× 21/2

d) MUx = 1/ x2

MUy = 1/ y2

MRS = (y/x)2 = 4